# Detecting Hidden Triggers: Mapping Non-Markov Reward Functions to Markov

### v2

@@ Abstract
Many Reinforcement Learning algorithms assume a Markov reward function to guarantee optimality. However, not all reward functions are known to be Markov. This paper proposes a framework for mapping non-Markov reward functions into equivalent Markov ones by learning a Reward Machine - a specialized reward automaton. Unlike the general practice of learning Reward Machines, we do not require a set of high-level propositional symbols from which to learn. Rather, we learn hidden triggers directly from data that encode them. Given their ability to model reward dependencies, we demonstrate the importance of learning Reward Machines versus their Deterministic Finite-State Automata counterparts. We formalize this distinction in our learning objective. Our mapping process is constructed as an Integer Linear Programming problem. We prove that our mappings provide consistent expectations for the underlying process. We empirically validate our approach by learning black-box non-Markov Reward functions in the Officeworld Domain. Additionally, we demonstrate the effectiveness of learning dependencies between rewards in a new domain, Breakfastworld.
@@ 1 Introduction
Reinforcement Learning (RL) is traditionally formulated as a sequential decision-making task modeled under a Markov Decision Process (MDP) [4]. This implies that both the reward and transition functions are Markov, depending solely on the current state and action. However, in complex environments, these functions are not always known and cannot always be assumed to be Markov. In fact, it has been shown that there are behaviors that cannot be encoded by any Markov reward function under certain state representations [1]. To address non-Markov dynamics, researchers often extend the state representation to enforce the Markov property, though this is usually at the cost of overcomplicating the state representation or making it excessively sparse. For example, Mnih et al. [19] stacked subsequent frames from the Atari games [3] to encode a suitable representation for training Deep Q-Networks. Similarly, deep, recurrent models have been used encode state histories for similar purposes [15, 11]. However, these sparse, deep models obscure the interpretation of rewards. To better understand them it is crucial to examine their dynamics under less sparse and more interpretable forms.
@@
A common strategy to represent non-Markov rewards in an interpretable format is to use the finite states of an automata to represent hidden or historical features that render the representation Markov. Historically, this has been accomplished with Deterministic Finite-State Automata (DFA) [5, 13, 12], but more recently, specialized reward automata, Reward Machines (RMs), have been used [32, 22, 23, 10]. We highlight two important distinctions between DFAs and RMs: 1.) DFAs are inherently less expressive than RMs as they only accept or reject traces. Consequently, when rewards are complex, multiple DFAs must be learned [13]. In contrast, RMs represent complex reward behavior in a single automaton. 2.) DFAs are only capable of representing patterns over $H=(S\times A)^{*}$, the history of states and actions. Conversely, RMs are capable of representing patterns over $H=(S\times A\times\mathbb{R})^{*}$, the history of states, actions, and rewards. This affords us a unique learning objective where rewards serve as memory allowing us to model reward dependencies.
@@
However, prior work for learning RMs assumes access to $P$, the set of high-level propositional symbols the RM operates over, and $L:S\times A\times S\rightarrow 2^{P}$, a labeling function mapping transitions to instantiations in $P$. This makes much of this prior work ungeneralizable. We instead infer $L$ by identifying minimal patterns in $H=(S\times A\times\mathbb{R})^{*}$, called hidden triggers, that differentiate reward outcomes. Rather than learn the RM directly, we map observed non-Markov rewards onto the cross-product of $U$, the state space of a candidate RM, and $S$, the observed state space, to form a Markov representation called the Abstract Reward Markov Decision Process (ARMDP). We infer our ARMDPs, and thus its decomposable RM, by solving an Integer Linear Program (ILP), lending itself to powerful off-the-shelf discrete optimization solvers.
Contributions
@@
We introduce a novel algorithm for mapping non-Markov reward functions to equivalent Markov ones. We show how leveraging $H=(S\times A\times\mathbb{R})^{*}$ in our learning objective can dramatically expedite learning in cases with multiple interdependent reward signals. We validate our approach in several ways. First, we demonstrate our ability to map non-Markov rewards sampled from black-box RMs in the Officeworld domain, originally introduced in the first RM paper [16]. Secondly, we evaluate the representative power of the ARMDP by extending DQN agents with ARMDP features to compare their learning profiles against various sparse Recurrent DQN (RDQN) variants. Finally, we validate our approach to represent complex reward function behavior in a new domain, Breakfastworld, featuring black-box RMs with interdependent reward signals. Importantly, we show how by leveraging $H=(S\times A\times\mathbb{R})^{*}$ we can expedite learning in these cases. Our contributions are three-fold:
@@
    We introduce a novel algorithm for RM learning without assumed access to $P$ and $L$,
@@
    We prove that the ARMDP is a suitable representation for maximizing reward expectations under non-Markov rewards,
@@ 2 Related Work
Decision-making with non-Markov reward functions is often represented under the Non-Markov Reward Decision Process (NMRDP) [27]. An NMRDP is similar to an MDP except that the reward function, $R:H\rightarrow\mathbb{R}$, maps state and action histories, $H=(S\times A)^{*}$, to reward values. However, as $H$ can become unwieldy in large state and action spaces, more succinct representations have been encoded using automata. The intuition of this approach lies in the fact that automata are finitely represented and offer a level of interpretation due to their structured form and symbolic representation.
Traditionally, the automata of choice has been the DFA, yielding promising results in identifying Markov representations for NMRDPs [5, 13, 12]. However, more recently, specialized reward automata, RMs, have been used to represent NMRDP behavior [16]. While DFAs operate over an input alphabet and emit true or false over traces, RMs extend this expressiveness by emitting reward values along each transition. RMs are particularly appealing as they offer succinct encoding of non-Markov reward behavior and task decomposition. They have also been extended by a broad range of RL algorithms that exploit RM structure to learn effective or optimal (in the discrete case) policies [7]. These include but are not limited to, specialized Q-Learning (QRM), Counterfactual experiences (CRM), Hierarchical Reinforcement Learning (HRM) and Reward Shaping (RSRM) for RMs [7, 17].
@@
A notable limitation of these prior works is that they assume access to a set of high-level symbols (the input language for the automata) and the relationship of the underlying state space to those symbols. In RM literature these are denoted $P$ and $L:S\times A\times S\rightarrow 2^{P}$ where $P$ is a set of symbols, separate from the NMRDP, and $L$ is a mapping of transitions from the NMRDP to instantiations in $P$, respectively. A similar assumption is often made for DFA learning, with exceptions being the work of Gaon and Brafman [13] as well as Christoffersen et al. [9] who learn symbolic representations from $H:(S\times A)^{*}$ to build their automata. Gaon and Brafman learn their DFA using L*, however, their approach was shown to be sample inefficient. Christoffersen et al. improved upon this by solving a discrete optimization problem for DFA learning, as outlined by Shvo et al. [24], which regularizes automata based on their size to better handle noise and enhance generalization. We improve upon these approaches by, instead, learning an RM without assumed access to $P$ and $L$. We show how by learning RMs we can model reward dependencies in a single automata by representing patterns in $H:(S\times A\times\mathbb{R})^{*}$.
@@ 3 Preliminaries
3.1 RL and Non-Markov Rewards
RL is a type of self-supervised learning in which an agent learns to interact with its environment to maximize its long-term reward expectation [25]. It is an iterative process in which the agent receives feedback in the form of rewards or punishments to adjust its behavior. RL is often represented as a sequential decision-making problem formalized using the MDP. An MDP is a 6-tuple $(S,A,T,R,\gamma,\rho)$, where $S$ is the state space, $A$ is the action space, $T:S\times A\times S\rightarrow[0,1]$ is a Markov probabilistic transition function, $R:S\times A\times S\rightarrow\mathbb{R}$ is a Markov reward function mapping states, actions and states to rewards, $\gamma\in[0,1)$ is the discount factor and $\rho:S\rightarrow[0,1]$ is the initial state distribution. We note that NMRDPs are similar to MDPs in that they share $(S,A,T,R,\rho)$, but $R:H\rightarrow\mathbb{R}$ in NMRDPs.
The solution for an MDP is a policy, $\pi:S\rightarrow A$, that maps states to actions. The value of a policy for any state, $s\in S$, at time point $t$ is the expected discounted return defined as follows:
$V_{\pi}(s)=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\mid s_{t}=s\right]$ (1)
The optimal policy denoted $\pi_{*}$, is determined by solving for the optimal state value function:
$V_{*}(s)=\max_{\pi}V^{\pi}(s)$ (2)
@@ 3.1 RL and Non-Markov Rewards
$Q_{*}(s,a)=\mathbb{E}\left[r_{t+1}+\gamma V_{*}(s_{t+1})\mid s_{t}=s,a_{t}=a\right]$ (3)
@@
where $q(s_{t},a_{t})$ is updated slowly according to $\alpha\in[0,1)$. When updating Equation (5) in an online fashion, typically an $\epsilon$-greedy strategy is employed where the agent follows a greedy policy but chooses a random action $\epsilon-$percent of the time.
Both strategies are guaranteed to converge to the optimal policy in the limit as each state-action pair is visited infinitely often, but only under the assumption that $T$ and $R$ are Markov. Otherwise, expectations have no stable solution.
@@ 3.2 Reward Machines
RMs are a specialized type of automata imposed over an underlying state and action space, $S$ and $A$ [16]. RMs operate over a set of propositional symbols, $P$, representing high-level events, separate from $S$. RMs are defined as a 5-tuple, $(U,u_{1},F,\delta_{u},\delta_{r})$, where $U$ is a finite state space, $u_{1}\in U$ is the initial state, $F$ is a set of terminal states where $F\cap U=\emptyset$, $\delta_{u}:U\times 2^{P}\rightarrow U\cup F$ is a transition function and $\delta_{r}:U\times 2^{P}\times U\cup F\rightarrow\mathbb{R}$ is a reward function. A labeling function, $L:S\times A\times S\rightarrow 2^{P}$, connects $S$ and $A$ to the RM by mapping transitions to instantiations in $P$.
@@
RMs are capable of modeling reward dependencies based on histories that can be represented as a regular language (e.g., loops, conditions, and interleaving). RMs were initially demonstrated on the Officeworld domain [16] shown in Figure 1 (a). Tasks (b-e) represent RMs sensitive to $P=\{c,m,o,*,A,B,C,D\}$ where $c$ is a coffee tile, $m$, the mail tile, $o$, the office tile, *, a plant tile, and $A,B,C,D$ are room IDs. $\neg$, $\lor$ and $\land$ are logical operators for negation, or, and and, respectively. The cross-product of $U$ and $S$ is Markov, which, in the discrete case, means we can solve these problems optimally.
@@ 4 Approach
From a set of observed, finite-length trajectories, $T_{o}$, sampled from an NMRDP, $(S,A,T,R)$, we learn a minimal RM $(U,u_{1},\delta_{u},\delta_{r})$ that renders rewards Markov to $S$, $A$ and $U$.
While $R$ in the NMRDP is unknown, we assume it can be sampled from and is representable as an RM. Thus, hereafter, we will refer to $R$ as a black-box RM. Our approach to solving this problem is first to learn an ARMDP, which is simply the MDP cross-product between an RM and an NMRDP, that enables the extraction of the associated RM. Learning of the ARMDP is formulated as an ILP. A minimal RM is an RM with the smallest hidden state space $|U|$, and the least number of transition edges between states.
The following sections are organized as follows. We begin by defining the ARMDP as the cross-product space between an RM and an NMRDP in Section 4.1. Section 4.2 describes a passive learning approach for learning ARMDPs from $T_{o}$, by solving an ILP. We later make this an active learning approach in Section 4.5 by incorporating a Q-Learning agent. Section 4.3 briefly outlines how to extract an RM from an ARMDP. Finally, Section 4.4 describes the theoretical insights of ARMDPs as a proxy for the NMRDP.
@@ 4.1 Abstract Reward Markov Decision Process
Given an NMRDP $(S,A,T,R)$, RM $(U,u_{1},\delta_{u},\delta_{r})$, and labeler, $L:S\times A\times S\rightarrow 2^{P}$, an ARMDP is an MDP $(\tilde{S},A,\tilde{T},\tilde{R},\gamma,\tilde{\rho})$ satisfying:
@@ 4.2 Formulation
We infer candidate ARMDPs by solving an ILP, that resolves reward conflicts over $T_{o}$. Let, $\tau_{m}\in T_{o}$, be the m-th trace in $T_{o}$ of length $l_{m}$. Further, let $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m},s_{n+1}^{m})$ be the n-th tuple of $\tau_{m}$. We define a reward conflict to be any two $\tau_{m,n}$ and $\tau_{\tilde{m},\tilde{n}}$ where $(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s_{\tilde{n}}^{\tilde{m}},a_{\tilde{n}}^{\tilde{m}},s_{\tilde{n}+1}^{\tilde{m}})$, but $r_{n+1}^{m}\neq r_{\tilde{n}+1}^{\tilde{m}}$.111 While we do not consider stochastic or noisy rewards in this work, we posit reward conflicts might be detected by sampling histories, in these cases. For example, if we define the history of $\tau_{m,n}$ to be $h(\tau_{m,n})=\{(s_{1}^{m},a_{1}^{m},r_{2}^{m},s_{2}^{m}),(s_{2}^{m},a_{2}^{m},r_{3}^{m},s_{3}^{m}),...,(s_{n-1}^{m},a_{n-1}^{m},r_{n}^{m},s_{n}^{m})\}$, then if $\tau_{m,n}$ is observed to conflict with $\tau_{\tilde{m},\tilde{n}}$ we might continually sample over $h(\tau_{m,n})$ and $h(\tau_{\tilde{m},\tilde{n}})$ to assess if their reward distributions are different in some statistical sense (e.g., difference of means). The motivation of our ILP is to map $T_{o}$ onto a valid ARMDP such that we resolve all reward conflicts observed in $T_{o}$. Intuitively, $\tilde{S}$, provides a degree of freedom over $U$, to do just that.
@@
Each $O_{m,n,i,j}\in O_{m,n}$ is a binary variable reflecting the decision to map $s_{n}^{m}$ to RM state $u_{i}\in U$ and $s_{n+1}^{m}$ to RM state $u_{j}\in U$. That is, if $O_{m,n,i,j}=1$, then $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m},s_{n+1}^{m})$ is mapped onto the ARMDP tuple, $((s_{n}^{m},u_{i}),a_{n}^{m},r_{n+1}^{m}(s_{n+1}^{m},u_{j}))$. We constrain each $O_{m,n}$ to choose a single mapping by enforcing that $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}$:
@@
We further constrain $O_{m,n}$ in instances where $n=1$ to satisfy the initial state requirement of Property 2 in the ARMDP. That is, $\forall m=1,2,...,M$:
@@
Next, we ensure the continuity of each trajectory in the mapping process. That is, we encode agreement between subsequent mappings of $O_{m,n}$ and $O_{m,n+1}$. To this end, we enforce the following constraint $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}-1,\forall j=1,2,...,K$:
$\sum_{i=1}^{K}O_{m,n,i,j}=\sum_{i=1}^{K}O_{m,n+1,j,i}$ (8)
Note that we swap indices $j$ and $i$ on the right-hand side. Equation (8) simply forces $s_{n+1}^{m}$ to be mapped to the same RM state in both $\tau_{m,n}$ and $\tau_{m,n+1}$. Eqs (6-8) form the basic scaffolding of our mapping.
@@
$I_{s,a,s^{\prime},i,j}$ is an indicator variable over the sum of all $O_{m,n,i,j}$ in $O$ with transition tuple, $(s,a,s^{\prime})$, and assignments $i,j\in\{1,2,...,K\}$. If $I_{s,a,s^{\prime},i,j}=1$, then our resulting ARMDP encodes the transition $((s,u_{i}),a,(s^{\prime},u_{j}))$ in $\tilde{T}$, and thus, $\delta_{u}(u_{j},L(s,a,s^{\prime}))=u_{j}$. If we let $(S,A,S)_{o}=\{(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}),\forall\tau_{m,n}\in\tau_{m},\forall\tau_{m}\in T_{o}\}$ be the set of all unique transitions in $T_{o}$, then to enforce the determinism of $\delta_{u}$ we impose that $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K$:
@@
Consider that for some $(s,a,s^{\prime})\in(S,A,S)_{o}$, and for some fixed $i\in\{1,2,...,K\}$ that when we constrain the sum $\forall j=1,2,...,K$ to be less than or equal to 1, we limit $\delta_{u}(u_{i},L(s,a,s^{\prime}))$ to have at most one outcome. We note that Equation (9) would be an equality when spanning $S\times A\times S$ instead of $(S,A,S)_{o}$. Equation (9), then, satisfies the determinism of $\delta_{u}$ and, therefore, Property 3 of the ARMDP.
@@
$I_{s,a,r,s^{\prime},i,j}$ is an indicator variable over the sum of all $O_{m,n,i,j}$ in $O$ with reward tuple, $(s,a,r,s^{\prime})$, and assignments $i,j\in\{1,2,...,K\}$. If $I_{s,a,r,s^{\prime},i,j}=1$, the ARMDP encodes $\tilde{R}((s,u_{i}),a,(s^{\prime},u_{j}))=r$, which is equivalent to $\delta_{r}(u_{i},L(s,a,s^{\prime}),u_{j})=r$ in the RM. If we let $r(s,a,s^{\prime})=\{r_{n}^{m},\forall\tau_{m,n}\in\tau_{m},\forall\tau_{m}\in T_{o}\mid(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s,a,s^{\prime})\}$ be the set of all observed rewards for some $(s,a,s^{\prime})$, then to enforce the determinism of $\delta_{r}$ we impose that $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K,\forall j=1,2,...,K$:

### v3

@@ Abstract
Many Reinforcement Learning algorithms assume a Markov reward function to guarantee optimality. However, not all reward functions are Markov. This paper proposes a framework for mapping non-Markov reward functions into equivalent Markov ones by learning specialized reward automata, Reward Machines. Unlike the general practice of learning Reward Machines, we do not require a set of high-level propositional symbols from which to learn. Rather, we learn hidden triggers, directly from data, that construct them. We demonstrate the importance of learning Reward Machines over their Deterministic Finite-State Automata counterparts given their ability to model reward dependencies. We formalize this distinction in our learning objective. Our mapping process is constructed as an Integer Linear Programming problem. We prove that our mappings form a suitable proxy for maximizing reward expectations. We empirically validate our approach by learning black-box, non-Markov reward functions in the Officeworld domain. Additionally, we demonstrate the effectiveness of learning reward dependencies in a new domain, Breakfastworld.
[…]
2377
@@ 1 Introduction
Reinforcement Learning (RL) is traditionally formulated as a sequential decision-making task modeled under a Markov Decision Process (MDP) [4]. This implies that both the reward and transition functions are Markov, depending solely on the current state and action. However, in complex environments, these functions are not always known and cannot always be assumed to be Markov. In fact, it has been shown that there are behaviors that cannot be encoded by any Markov reward function under certain state representations [1]. To address non-Markov dynamics, researchers often extend the state representation to enforce the Markov property, though this is usually at the cost of overcomplicating the state representation or making it excessively sparse. For example, Mnih et al. [18] stacked subsequent frames from the Atari games [3] to encode a suitable representation for training Deep Q-Networks (DQNs). Similarly, Recurrent DQNs (RDQNs) have been used to encode state histories for similar purposes [14]. However, these sparse, deep models obscure the interpretation of rewards. To better understand them, it is crucial to examine their dynamics under less sparse and more interpretable forms.
@@
To this end, automata have been used to compactly represent the hidden or historical features of non-Markov rewards. While Deterministic Finite-State Automata (DFAs) have traditionally served this role [5, 12, 11], more recently, specialized reward automata, Reward Machines (RMs), have been used [31, 21, 22, 10]. We highlight two distinctions between DFAs and RMs: 1.) DFAs are inherently less expressive than RMs as they only accept or reject traces. Consequently, when rewards are complex, multiple DFAs must be learned [12]. In contrast, RMs can encode complex reward behavior in a single automaton. 2.) DFAs are only capable of representing patterns over $H=(S\times A)^{*}$, the history of states and actions. Conversely, RMs can represent patterns over $H=(S\times A\times\mathbb{R})^{*}$, which include rewards. This affords us a unique learning objective where rewards serve as memory allowing us to model their dependencies.
@@
However, prior work on learning RMs assumes access to $P$, a set of high-level propositions, and $L:S\times A\times S\rightarrow 2^{P}$, a labeling function mapping transitions to instantiations in $P$. This makes much of this prior work ungeneralizable. We, instead, infer $L$ by identifying minimal patterns in $H=(S\times A\times\mathbb{R})^{*}$, called hidden triggers, that differentiate reward outcomes. Rather than learn RMs directly, we map observed non-Markov rewards onto the cross-product of $U$, the state space of a candidate RM, and $S$, the observed state space, to form a Markov representation called the Abstract Reward Markov Decision Process (ARMDP). We infer an ARMDP, and thus its decomposable RM, by solving an Integer Linear Program (ILP), lending itself to powerful off-the-shelf discrete optimization solvers.
@@
Contributions: We introduce a novel algorithm for mapping non-Markov reward functions to equivalent Markov ones. We show how by leveraging $H=(S\times A\times\mathbb{R})^{*}$ in our learning objective we can dramatically expedite learning in cases with interdependent reward signals. We validate our approach by demonstrating our ability to learn black-box RMs from the Officeworld domain, originally introduced in the first RM paper [15]. We then evaluate the representative power of the ARMDP by extending DQN agents with ARMDP features to compare their learning profiles against various sparse RDQN variants. Finally, we validate our approach to represent complex reward function behavior in a new domain, Breakfastworld, featuring black-box RMs with interdependent reward signals. We show how by leveraging $H=(S\times A\times\mathbb{R})^{*}$ in these experiments we can expedite learning. Our main contributions are three-fold:
@@
    We introduce a novel algorithm for learning RMs without assumed access to $P$ and $L$,
@@
    We prove that the ARMDP is a suitable proxy for maximizing reward expectations under non-Markov rewards,
@@ 2 Related Work
Decision-making with non-Markov reward functions is often represented as a Non-Markov Reward Decision Process (NMRDP) [26]. An NMRDP is similar to an MDP except that the reward function, $R:H\rightarrow\mathbb{R}$, maps state and action histories, $H=(S\times A)^{*}$, to reward values. However, as $H$ can become unwieldy due to the curse of history, more succinct representations have been encoded using automata. The intuition of this approach lies in the fact that automata are finitely represented and offer a level of interpretation due to their structured form and symbolic representation.
Traditionally, the automata of choice has been the DFA, yielding promising results in identifying Markov representations for NMRDPs [5, 12, 11]. More recently, specialized reward automata, RMs, have also been used [15]. While DFAs operate over an input alphabet and emit true or false over traces, RMs extend this expressiveness by emitting reward values along each transition. RMs are particularly appealing as they offer succinct encoding of non-Markov reward behavior and task decomposition. They have also been extended by a broad range of RL algorithms that exploit RM structure to learn effective or optimal (in the discrete case) policies [7]. These include but are not limited to, specialized Q-Learning (QRM), Counterfactual experiences (CRM), Hierarchical Reinforcement Learning (HRM) and Reward Shaping (RSRM) for RMs [7, 16].
@@
A notable limitation of these prior works is that they assume access to a set of high-level symbols (the input language for the automata) and the relationship of the underlying state space to those symbols. In RM literature, these are denoted $P$ and $L:S\times A\times S\rightarrow 2^{P}$ where $P$ is a set of symbols, separate from the NMRDP, and $L$ is a mapping of transitions from the NMRDP to instantiations in $P$, respectively. A similar assumption is often made for DFA learning, with exceptions being the work of Gaon and Brafman [12] as well as Christoffersen et al. [9] who learn symbolic representations from $H=(S\times A)^{*}$ to build their automata. Gaon and Brafman learn their DFA by applying L*, however, their approach was shown to be sample inefficient. Christoffersen et al. improved upon this by solving a discrete optimization problem for DFA learning, as outlined by Shvo et al. [23], which regularizes automata based on their size to better handle noise and enhance generalization. We improve upon these approaches by, instead, learning an RM without assumed access to $P$ and $L$. We show how, by learning RMs, we can model reward dependencies in a single automata by representing patterns in $H=(S\times A\times\mathbb{R})^{*}$.
@@ 3 Preliminaries
3.1 Reinforcement Learning and Non-Markov Reward
RL is a type of self-supervised learning in which an agent learns to interact with its environment to maximize its long-term reward expectation [24]. It is an iterative process in which the agent receives feedback in the form of rewards or punishments to adjust its behavior. RL is formalized using the MDP. An MDP is a 6-tuple $(S,A,T,R,\gamma,\rho)$, where $S$ is the state space, $A$ is the action space, $T:S\times A\times S\rightarrow[0,1]$ is a Markov probabilistic transition function, $R:S\times A\times S\rightarrow\mathbb{R}$ is a Markov reward function mapping states, actions and states to rewards, $\gamma\in[0,1)$ is the discount factor and $\rho:S\rightarrow[0,1]$ is the initial state distribution. Briefly, NMRDPs are similar to MDPs in that they share $(S,A,T,R,\gamma,\rho)$; however, $R:H\rightarrow\mathbb{R}$ maps histories of states and actions to rewards, instead.
The solution for an MDP is a policy, $\pi:S\rightarrow A$, that maps states to actions. The value of a policy for any state, $s\in S$, at time point $t$ is the expected discounted return, defined as follows:
$V_{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\mid s_{t}=s\right]$ (1)
where $\mathbb{E}_{\pi}$ is the expected reward from starting in $s$ and following $\pi$. The optimal policy, denoted $\pi_{*}$, is determined by solving for the optimal state value function:
$V_{*}(s)=\max_{\pi}V_{\pi}(s)$ (2)
@@ 3.1 Reinforcement Learning and Non-Markov Reward
$Q_{*}(s,a)=\mathbb{E}_{\pi}\left[r_{t+1}+\gamma V_{*}(s_{t+1})\mid s_{t}=s,a_{t}=a\right]$ (3)
@@
where $q(s_{t},a_{t})$ is updated slowly according to $\alpha\in[0,1)$. When updating Equation (5) in an online fashion, typically an $\epsilon$-greedy strategy is employed where the agent follows a greedy policy but chooses a random action $\epsilon-$percent of the time. Both strategies are guaranteed to converge to the optimal policy in the limit as each state-action pair is visited infinitely often, but only under the assumption that $T$ and $R$ are Markov. Otherwise, expectations have no stable solution.
@@ 3.2 Reward Machines
RMs are a specialized type of automata imposed over an underlying state and action space, $S$ and $A$ [15]. RMs operate over a set of propositional symbols, $P$, representing high-level events, separate from $S$. RMs are defined as a 5-tuple, $(U,u_{1},F,\delta_{u},\delta_{r})$, where $U$ is a finite state space, $u_{1}\in U$ is the initial state, $F$ is a set of terminal states where $F\cap U=\emptyset$, $\delta_{u}:U\times 2^{P}\rightarrow U\cup F$ is a transition function and $\delta_{r}:U\times 2^{P}\times(U\cup F)\rightarrow\mathbb{R}$ is a reward function. A labeling function, $L:S\times A\times S\rightarrow 2^{P}$, connects $S$ and $A$ to the RM by mapping transitions to instantiations in $P$.
@@
RMs are capable of modeling reward dependencies based on histories that can be represented as a regular language (e.g., loops, conditions, and interleaving). RMs were initially demonstrated on the Officeworld domain [15] shown in Figure 1 (a). Tasks (b-e) represent RMs sensitive to $P=\{c,m,o,*,A,B,C,D\}$ where $c$ is a coffee tile, $m$, the mail tile, $o$, the office tile, *, a plant tile, and $A,B,C,D$ are room IDs. $\neg$, $\lor$ and $\land$ are logical operators for negation, or, and and, respectively. The cross-product of an NMRDP and an RM can be used to form a Markov representation for predicting both $R$ and $T$.
@@ 4 Approach
Given a finite set of finite-length trajectories, $T_{o}$, sampled from an NMRDP, $(S,A,T,R,\gamma,\rho)$, we learn a minimal RM $(U,u_{1},\delta_{u},\delta_{r})$ that renders rewards Markov to $S$, $A$ and $U$.
While $R$ in the NMRDP is unknown, we assume it can be sampled from and is representable as an RM. Hereafter, we will refer to $R$ as a black-box RM. Our approach for solving this problem is to learn an ARMDP, which is simply the MDP cross-product between an RM and an NMRDP that enables the extraction of the associated RM. Learning of the ARMDP is formulated as an ILP. A minimal RM is an RM with the smallest hidden state space $|U|$, and the least number of transition edges between different RM states.
The following sections are organized as follows: We begin by defining the ARMDP as the cross-product space between an RM and an NMRDP in Section 4.1. Section 4.2 describes a passive learning approach for learning ARMDPs from $T_{o}$, by solving an ILP. We later make this an active learning approach in Section 4.5 by incorporating a Q-Learning agent. Section 4.3 briefly outlines how to extract an RM from an ARMDP. Finally, Section 4.4 describes the theoretical insights of ARMDPs as a proxy for the NMRDP.
@@ 4.1 Abstract Reward Markov Decision Process
Given an NMRDP $(S,A,T,R,\gamma,\rho)$, RM $(U,u_{1},\delta_{u},\delta_{r})$, and labeler, $L:S\times A\times S\rightarrow 2^{P}$, an ARMDP is an MDP, $(\tilde{S},A,\tilde{T},\tilde{R},\gamma,\tilde{\rho})$, satisfying:
@@ 4.2 Formulation
We infer candidate ARMDPs by solving an ILP that resolves reward conflicts over $T_{o}$. Let $\tau_{m}\in T_{o}$ be the m-th trace in $T_{o}$ of length $l_{m}$. Further, let $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m},s_{n+1}^{m})$ be the n-th tuple of $\tau_{m}$. We define a reward conflict to be any two $\tau_{m,n}$ and $\tau_{\tilde{m},\tilde{n}}$ where $(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s_{\tilde{n}}^{\tilde{m}},a_{\tilde{n}}^{\tilde{m}},s_{\tilde{n}+1}^{\tilde{m}})$, but $r_{n+1}^{m}\neq r_{\tilde{n}+1}^{\tilde{m}}$. The motivation of our ILP is to map $T_{o}$ onto a valid ARMDP such that we resolve all reward conflicts observed in $T_{o}$. Intuitively, $\tilde{S}$ provides a degree of freedom over $U$ to do just that.
@@
Each $O_{m,n,i,j}\in O_{m,n}$ is a binary variable reflecting the decision to map $s_{n}^{m}$ to RM state $u_{i}\in U$ and $s_{n+1}^{m}$ to RM state $u_{j}\in U$. That is, if $O_{m,n,i,j}=1$, then $\tau_{m,n}=(s_{n}^{m},a_{n}^{m},r_{n+1}^{m},s_{n+1}^{m})$ is mapped onto the corresponding ARMDP tuple of the form, $((s_{n}^{m},u_{i}),a_{n}^{m},r_{n+1}^{m}(s_{n+1}^{m},u_{j}))$. As such, each $\tau_{m,n}$ introduces $K^{2}$ variables through its respective $O_{m,n}$ matrix. We constrain each $O_{m,n}$ to choose a single mapping by enforcing that $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}$:
@@
We further constrain $O_{m,n}$ where $n=1$ to satisfy the initial state requirement of Property 2 in the ARMDP. That is, $\forall m=1,2,...,M$:
@@
Next, we ensure the continuity of each trajectory in the mapping process. That is, we encode agreement between subsequent mappings of all $O_{m,n}$ and $O_{m,n+1}$ pairs by enforcing the following constraint $\forall m=1,2,...,M,\forall n=1,2,...,l_{m}-1,\forall j=1,2,...,K$:
$\sum_{i=1}^{K}O_{m,n,i,j}=\sum_{j^{\prime}=1}^{K}O_{m,n+1,j,j^{\prime}}$ (8)
Equation (8) simply forces $s_{n+1}^{m}$ to be mapped to the same RM state in both $\tau_{m,n}$ and $\tau_{m,n+1}$. Equations (6-8) form the scaffolding of our mapping.
@@
$I_{s,a,s^{\prime},i,j}$ is an indicator variable representing the sum of all $O_{m,n,i,j}$ in $O$ sharing the transition tuple $(s,a,s^{\prime})$ with assignments $i,j\in{1,2,...,K}$. When $I_{s,a,s^{\prime},i,j}=1$, it acts as a hidden trigger that highlights the sensitivity of $L$ to $(s,a,s^{\prime})$ and enforces the transition $\delta_{u}(u_{j},L(s,a,s^{\prime}))=u_{j}$, in the RM, thus governing $\tilde{T}$. We later consider the sensitivity of hidden triggers to rewards in Property 4.
Let $(S,A,S)_{o}=\{(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\mid m\in\{1,2,...,M\},n\in\{1,2,...,l_{m}\}\}$ be the set of all unique transitions in $T_{o}$. Then to enforce the determinism of $\delta_{u}$, we impose that $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K$:
@@
Consider that for any $(s,a,s^{\prime})\in(S,A,S)_{o}$ and for any $i\in\{1,2,...,K\}$ that when we constrain the sum $\forall j=1,2,...,K$, in Equation (9), to be less than or equal to 1, we limit $\delta_{u}(u_{i},L(s,a,s^{\prime}))$ to have at most one outcome. With this, Equation (9) satisfies the determinism of $\delta_{u}$ and Property 3 of the ARMDP.
@@
$I_{s,a,r,s^{\prime},i,j}$ is an indicator variable representing the sum over all $O_{m,n,i,j}$ in $O$ sharing the reward tuple $(s,a,r,s^{\prime})$ with assignments $i,j\in\{1,2,...,K\}$. When $I_{s,a,r,s^{\prime},i,j}=1$, it imposes the following reward emission, $\delta_{r}(u_{i},L(s,a,s^{\prime}),u_{j})=r$, in the RM.
Let $r(s,a,s^{\prime})=\{r_{n+1}^{m}\mid m\in\{1,2,...,M\},\hskip 2.0ptn\in\{1,2,...,l_{m}\},\\ (s_{n}^{m},a_{n}^{m},s_{n+1}^{m})=(s,a,s^{\prime})\}$ be the set of all observed rewards for some $(s,a,s^{\prime})$. Then to enforce the determinism of $\delta_{r}$, we impose that $\forall(s,a,s^{\prime})\in(S,A,S)_{o},\forall i=1,2,...,K,\forall j=1,2,...,K$: